# A. CamillettiAI weather from Euro-Atlantic regimes

### v1

@@ 3.3 Target fields: temperature and precipitation anomalies
The daily two-meter temperature and total precipitation anomalies are computed following the same procedure at points 1 to 6 described in Sec. 3.1. Following the same pre-processing, the relevant relations between WR and target anomalies are preserved, allowing the AI model to exploit them for the reconstruction. The daily variables are then aggregated monthly: the daily mean two-meter temperature anomaly is averaged every day in a month while the daily total precipitation anomaly is summed. The monthly average two-meter temperature and total precipitation anomalies are the target variables that the AI model is asked to reconstruct.
@@ 3.4 SEAS5 bias-corrected forecast
Despite the AI model is trained on the $I_{\rm wr}$ computed from ERA5 $Z500$ standardized anomalies, $\tilde{\Phi}(r,\lambda,\phi)$, it is interesting to evaluate its performance in reconstructing the monthly average two-meter temperature and total precipitation anomalies using the predicted $I_{\rm wr}$ computed from the bias-corrected SEAS5 $Z500$ forecast, $I_{\rm wr}^{\rm SEAS5}$. To obtain $I_{\rm wr}^{\rm SEAS5}$ we project the forecast of the bias-corrected monthly mean $Z500$ standardized anomalies for each ensemble member $n$ and lead-time month $l$, $\tilde{\Phi}^{\rm SEAS5}(t,\lambda,\phi,l,n)$, to the seven mean cluster centers computed from the ERA5 daily anomalies, $\tilde{\Phi}_{\rm wr}(\lambda,\phi)$, (see Sec. 3.1). The calculations follow closely Eqs. 1 and 2, with the additional dimensions $l$ and $n$.
@@ 4 AI-model
Figure 2: Schematic representation of the model. The input vector $x$ with the $n$ WR indices, the year, and the encoded month is passed as an input to the model. After each layer, the output $y$ is passed to a ReLU activation function $\max(0,y)$. The $N=28$ dimensional vector, output of the second linear ResNet block, is stacked pixel-by-pixel onto the static data (latitude and longitude grids, land-sea mask, and digital terrain model). The resulting $(N+4)\times n_{\lambda}\times n_{\phi}$ tensor, where $n_{\phi}$ and $n_{\lambda}$ are the number of latitude and longitude grid points, respectively, is the input of the first convolutional ResNet layer. The number of features is then reduced by two by each of the next convolutional ResNet blocks.
@@
The AI model is designed to reconstruct the monthly average of the standardized two-meter temperature and total precipitation anomalies in Europe (Sec. 3.3) from the monthly average WR index, $I_{\rm wr}$ (Sec. 3.1).
4.1 Model architecture and training
Fig. 2 schematically describes the model architecture. The monthly average $I_{\rm wr}$, for each WR, is passed along with the corresponding encoded month and year as input to the model. The month is encoded in twelve categories before all the inputs are processed first by a linear layer and subsequently to a sequence of two ResNet blocks, which takes the 20-dimensional input (seven WR, the year, and the twelve-dimensional encoded month) and yields a vector of $N=28$ dimensions. The $N$-dimensional vector is stacked pixel-by-pixel with the two-dimensional static data: the latitude grid, the longitude grid, the land-sea mask, and the digital terrain model (DTM) of the European region with the same resolution as the output two-dimensional field. The resulting $(N+4)\times n_{\lambda}\times n_{\phi}$ tensor, where $n_{\phi}$ and $n_{\lambda}$ are the numbers of latitude and longitude grid points, respectively, is the input of the next convolutional ResNet blocks, which reduce the number of features to one, i.e., to a two-dimensional field representing the target anomalies.
The AI model is trained separately to reconstruct two-meter temperature and total precipitation anomalies. It learns to approximate the target anomalies by adjusting its internal $\approx~{}30$k parameters to minimize the mean squared error (MSE) between the predicted and observed anomaly fields. To properly assess model performance and its generalization ability, the dataset is divided into distinct training and test periods. The training set spans from January 1940 to December 2010 , including the climatological period used to compute anomalies and define the WR indices. The model’s performance can be slightly influenced by the order in which the input and output pairs of the test dataset are passed to the model. We use random ordering and repeat the training procedure six times to remove better the dependence of the model’s performance from the training data random ordering. The test set covers the months from January 2011 to December 2024 , and a portion of this period, from January 2011 to December 2018, is also used for validation. Training is halted when the validation MSE reaches a minimum to prevent overfitting. It is important to note that the evaluation period includes the validation dataset used during model training. While this could potentially introduce bias into the reconstruction skill estimates, this issue is addressed and quantitatively assessed in Appendix D, where we demonstrate that any resulting bias is negligible. We choose to present results over the extended test period to ensure statistical robustness and enhance the reliability of the evaluation through a larger sample size. The limited size of the training dataset (70 years) constrains the choice of model architecture, preventing the use of larger and more complex models. Such models, with their higher number of parameters, would not only be computationally inefficient but also more prone to overfitting in data-scarce settings.
@@ 4.2 Evaluation metrics
Figure 3: AI model spatial MAE (top), ACC (middle) and CE (bottom) in reconstructing the winters (first and third columns) and summers (second and fourth columns) mean two-meter temperature and total precipitation anomalies using seven WR indices in the test years (from January 2011 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC (CE) higher than 0.5 (0.3).
@@
${\rm{MAE}}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|~{}.$ (3)
@@
is a statistical measure to assess the similarity between predicted and observed anomalies in a dataset. It is commonly applied in climate science to evaluate deterministic forecast accuracy. ACC quantifies how well the anomalies correlate between predictions and observations. Given a set of anomalies for $n$ observations, where $y_{i}$ represents the observed anomalies and $\hat{y}_{i}$ the predicted ones, ACC is computed as
${\rm{ACC}}=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})\left(\hat{y}_{i}-\bar{\hat{y}}\right)}{\sqrt{\sum_{i=1}^{n}\left(y_{i}-\bar{y}\right)^{2}}\sqrt{\sum_{i=1}^{n}\left(\hat{y}_{i}-\bar{\hat{y}}\right)^{2}}}~{},$ (4)
where $\bar{y}$ and $\bar{\hat{y}}$ denote the mean values of observed and predicted anomalies, respectively. The ACC ranges from $-1$ to $1$, with values close to $1$ indicating a strong positive correlation, meaning that the model effectively captures the variations in anomalies. Conversely, values near 00 suggest no correlation, while negative values indicate an inverse relationship between predicted and observed anomalies.
To assess whether an observed correlation is statistically significant, we compute the associated p-value. The p-value represents the probability of obtaining higher or equal correlation assuming that there is no true correlation (i.e., under the null hypothesis of zero correlation). In this context, a p-value below 0.05 is typically considered statistically significant, indicating that there is less than a 5% probability that the observed correlation occurred by chance. Thus, correlations with $p<0.05$ can be interpreted as significant at the 95% confidence level.
@@
is a statistical measure used to evaluate the predictive skill of a model by comparing its performance to that of a validation climatology [Bürger, 2007], thus indicating the shared variance between the observed and estimated data throughout the verification period. Given $n$ observations, where $y_{i}$ represents the actual values, $\hat{y}_{i}$ the predicted values, and $\bar{y}$ the mean of observed values, the CE is computed for each grid point as,
${\rm{CE}}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}~{}.$ (5)
A value of $\rm{CE}=1$ indicates a perfect prediction, while $\rm{CE}=0$ suggests that the model performs no better than the validation climatology. Negative values imply that the model performs worse than the reference. In this study we use the climatology of the 2011 – 2024 testing period as validation climatology.
@@
To evaluate the aforementioned deterministic metrics on SEAS5, if no otherwise specified, we always use the ensemble mean as the predicted value.
@@ 5 Results
Figure 4: 2011 – 2024 winters (first and third columns) and summers (second and fourth columns) MAE (top row), ACC (middle row) and CE (bottom row) of the two-meter temperature anomalies and precipitation anomalies, reconstructed by the AI model trained with seven WR (red), four WR (orange), NAO (light-blue) and zero (blue) indices.
In the following sections, we evaluate the performance of the AI model in reconstructing monthly mean two-meter temperature and total precipitation anomalies in winter (DJF) and summer (JJA) from January 2011 to December 2024 . Sec. 5.1 presents the spatial reconstruction of these anomalies by the AI model trained with the seven year-round WR indices. We further investigate the model’s skill with respect to: (a) the number of WR indices used for training and as input (Sec. 5.2); (b) the impact of errors on $I_{\rm wr}$ on the reconstruction (Sec. 5.3); and (c) the use of $I_{\rm wr}^{\rm SEAS5}$ as input (Sec. 5.4). The first two aspects address how the reconstruction quality degrades, identifying a threshold to achieve a better seasonal forecast than SEAS5. The third point examines the accuracy of the reconstructed anomalies when using WR indices predicted by SEAS5. It can be interpreted as a hybrid NWP-AI framework for forecasting European mean two-meter temperature and total precipitation in winter and summer.
Although the AI model reconstructs anomalies over the entire European region as defined in Sec. 3.3, in the following analysis we restrict the evaluation to land areas only.
@@ 5.1 Anomaly reconstruction with seven WR
Fig. 3 shows the spatial MAE, ACC and CE in reconstructing the European monthly mean two-meter temperature and total precipitation anomaly using the seven-year-round WR indices, $I_{\rm wr}$, in the winters and summers during the test years.
@@
For the two-meter temperature anomaly, the MAE maps indicate low reconstruction errors across most of Europe, especially in central and southern regions, with values frequently below 0.5. Errors are slightly higher in northern and eastern areas, particularly during winter, which is related to the higher variability and amplitude of the two-meter temperature anomalies in those regions. The ACC maps show high correlations (generally above 0.5) between reconstructed and observed anomalies, particularly over western and central Europe in both winter and northern east Europe in summer, but also in some part of the Mediterranean region. This indicates the model effectively captures temporal variability in these areas. The CE metric partially mirrors this pattern, with values approaching 0.3 in some areas, signifying that the model performs appreciably better than a climatological reference.
@@
In the case of precipitation, the MAE values are more spatially variable and generally higher than for temperature, which reflects the inherently more chaotic and less predictable nature of precipitation. Nonetheless, several regions—especially parts of western Europe—exhibit relatively low MAE, particularly in winter. The ACC values for precipitation are moderate (typically between 0.4 and 0.7), with some improvement in winter relative to summer, suggesting that the WR-based model can partially capture precipitation variability, albeit with reduced skill compared to temperature. CE values for precipitation exhibit a similar spatial pattern, with positive skill in central Europe during winter but a marked decline in some southern and eastern regions during summer.
@@
Summer extratropical variability is less influenced by the large amplitude stationary waves and tropical forcing that characterize the boreal winter climate. As a result, the summer weather in the Northern Hemisphere is more influenced by regional-scale processes, such as mesoscale convective weather systems and land-atmosphere coupling [Ambrizzi et al., 1995, Koster et al., 2000, Schubert et al., 2011]. This explains the lower skill reached by the AI model in summer compared to winter in all the skills tested. In particular, summer’s CE show that the model using the Euro-Atlantic WR is unable to appropriately reconstruct the mean two-meter temperature and total precipitation anomalies in the Mediterranean and Black Sea regions. For these areas, the reconstruction could be improved by including the Mediterranean large-scale circulation [Mastrantonas et al., 2021] and adding other predictors such as soil moisture or rainfall deficit indices [Vautard et al., 2007].

### v2

@@ 3.3 Target fields: temperature and precipitation anomalies
The daily two-meter temperature and total precipitation anomalies are computed following the same procedure at points 1 to 3 described in Sec. 3.1. Following the same pre-processing, the relevant relations between WR and target anomalies are preserved, allowing the AI models to exploit them for the reconstruction. Note that we do not normalize the anomalies by the calendar standard deviation, so we express the anomalies in Kelvin (K) for the monthly mean two-meter temperature and in centimeters (cm) for the monthly total precipitation.
The daily variables are then aggregated monthly: the daily mean two-meter temperature anomaly is averaged over every day in the month, while the daily total precipitation anomaly is summed. The monthly average two-meter temperature and total precipitation anomalies are the target variables that the AI models are asked to reconstruct.
@@ 3.4 SEAS5 bias-corrected forecast
Despite the AI models are trained on the $I_{\rm wr}$ computed from ERA5 $Z500$ standardized anomalies, $\tilde{\Phi}(r,\lambda,\phi)$, it is interesting to evaluate its performance in reconstructing the monthly average two-meter temperature and total precipitation anomalies using the predicted $I_{\rm wr}$ computed from the bias-corrected SEAS5 $Z500$ forecast, $I_{\rm wr}^{\rm SEAS5}$. To obtain $I_{\rm wr}^{\rm SEAS5}$ we project the forecast of the bias-corrected monthly mean $Z500$ standardized anomalies for each ensemble member $m$ and lead-time month $l$, $\tilde{\Phi}^{\rm SEAS5}(t,\lambda,\phi,l,n)$, to the seven mean cluster centers computed from the ERA5 daily anomalies, $\tilde{\Phi}_{\rm wr}(\lambda,\phi)$, (see Sec. 3.1). The calculations follow closely Eqs. 1 and 2, with the additional dimensions $l$ and $m$.
@@ 4 AI-model
Figure 2: Schematic representation of the model’s architecture. The input vector $x$, which includes the $n$ WR indices, the year, and the encoded month, is passed as input to the model. After each layer, the output $y$ is passed to a ReLU activation function $\max(0,y)$. The $N=28$ dimensional vector, output of the second linear ResNet block, is stacked pixel-by-pixel onto the static data (latitude and longitude grids, land-sea mask, and digital terrain model). The resulting $(N+4)\times n_{\lambda}\times n_{\phi}$ tensor, where $n_{\phi}$ and $n_{\lambda}$ are the number of latitude and longitude grid points, respectively, is the input of the first convolutional ResNet layer. The number of features is then reduced by two by the second ResNet block and to one (temperature or precipitation field) by the final ResNet layer
@@
The AI models are designed to reconstruct the monthly average of the standardized two-meter temperature and total precipitation anomalies in Europe (Sec. 3.3) from the monthly average WR index, $I_{\rm wr}$ (Sec. 3.1).
4.1 Models’ architecture and training
Fig. 2 schematically describes the models’ architecture. The monthly average $I_{\rm wr}$, for each WR, is passed along with the corresponding encoded month and year as input to the models. The month is encoded in twelve categories using one-hot encoding. One-hot encoding is a widely used technique in machine learning and data processing for representing categorical variables as numerical vectors. In our scenario, we encode the month = 1,…,12, into a twelve-dimensional vector, e.g., January (month = 1) is encoded in (1,0,…,0); February (month = 2) is encoded in (0,1,0, …,0). The inputs are then processed first by a linear layer and subsequently by a sequence of two ResNet blocks. ResNet (Residual Network) introduces shortcut connections, allowing the model to learn residual mappings instead of direct transformations, which makes it easier to train deep architectures. The use of ResNet blocks helps mitigate the vanishing gradient problem, ensuring stable and efficient training of deeper networks, which is particularly beneficial when working with limited data. The ResNet blocks take the 20-dimensional input (seven WR, the year, and the twelve-dimensional encoded month) and yield a vector of $N=28$ dimensions. Expanding the feature space is a common strategy in NN, as it enables the model to better capture and process the information in the input data by learning a more expressive representation of the temporal predictors. The $N$-dimensional vector is stacked pixel-by-pixel with the two-dimensional static data: the latitude grid, the longitude grid, the land-sea mask, and the digital terrain model (DTM) of the European region with the same resolution as the output two-dimensional field. Stacking these features pixel-by-pixel with the static spatial data enables the model to condition its reconstructions on both large-scale atmospheric drivers and local geographical features. The resulting $(N+4)\times n_{\lambda}\times n_{\phi}$ tensor, where $n_{\phi}$ and $n_{\lambda}$ are the numbers of latitude and longitude grid points, respectively, is the input of the following convolutional ResNet blocks, which reduce the number of features to one, i.e., to a two-dimensional field representing the target anomalies. Convolutional layers capture both local spatial patterns and cross-grid dependencies, enabling the model to learn large-scale drivers and their geographically-modulated regional responses. We use 3x3 kernels in all the convolutional layers. These layers combine the information encoded in the feature vector with the static spatial data by selecting the components of the feature vector most relevant to each grid point. A 3×3 kernel is sufficiently small to capture local structure. We noticed that increasing the number of convolutional ResNet layers did not improve the performance of the model. We employ ReLU activation functions to introduce non-linearity in a computationally efficient way, enabling the model to capture complex patterns.
The AI models are trained separately to reconstruct two-meter temperature and total precipitation anomalies. They learn to approximate the target anomalies by adjusting their internal $\approx 30$k parameters to minimize the training loss between the predicted and observed anomaly fields. We choose the mean squared error (MSE) as training loss, defined in Sec. 4.2. For more details about the training setup and hyperparameters, the reader can refer to Appendix C. To properly assess models’ performance and their generalization ability, the dataset is divided into distinct training, validation, and test periods. The training set spans from January 1940 to December 2005 . The months from January 2006 to December 2010 are used to avoid overfitting, by halting the training when the validation MSE reaches a minimum. To finally assess the performance of the model, the remaining months from January 2011 to December 2024 are used. Note that the climatology period, used to compute the anomalies of the WR, does not include the test dataset.
The models’ performance can be slightly influenced by the order in which the input and output pairs of the train dataset are passed to the models during training. We use random ordering and repeat the training procedure six times to account for the dependence of the model’s performance on the training data random ordering.
The limited size of the training dataset (65 years) constrains the choice of models’ architecture, preventing the use of larger and more complex architectures. Such architectures, with their higher number of parameters, would not only be computationally inefficient but also more prone to overfitting in data-scarce settings.
@@ 4.2 Evaluation metrics
${\mathrm{MAE}}=\frac{1}{n}\sum_{i=1}^{n}\left|y_{i}-\hat{y}_{i}\right|\penalty 10000\ .$ (3)
@@
ACC is a statistical measure to assess the similarity between predicted and observed anomalies in a dataset. It is commonly applied in climate science to evaluate the accuracy of deterministic forecasts. ACC quantifies how well the anomalies correlate between predictions and observations. Given a set of anomalies for $n$ observations, where $y_{i}$ represents the observed anomalies and $\hat{y}_{i}$ the predicted ones, anomaly correlation coefficient (ACC) is computed as
${\mathrm{ACC}}=\frac{\sum_{i=1}^{n}(y_{i}-\bar{y})\left(\hat{y}_{i}-\bar{\hat{y}}\right)}{\sqrt{\sum_{i=1}^{n}\left(y_{i}-\bar{y}\right)^{2}}\sqrt{\sum_{i=1}^{n}\left(\hat{y}_{i}-\bar{\hat{y}}\right)^{2}}}\penalty 10000\ ,$ (4)
where $\bar{y}$ and $\bar{\hat{y}}$ denote the mean values of observed and predicted anomalies, respectively. The ACC ranges from 1-1 to $1$, with values close to $1$ indicating a strong positive correlation, meaning that the model effectively captures the variations in anomalies. Conversely, values near 0 suggest no correlation, while negative values indicate an inverse relationship between predicted and observed anomalies.
To assess whether an observed correlation is statistically significant, we compute the associated two-tailed p-value, derived from the exact distribution of the sample correlation coefficient, which follows a scaled beta distribution. The two-tailed p-value represents the probability of obtaining a higher or equal correlation assuming that there is no actual correlation (i.e., under the null hypothesis of zero correlation). In this context, a p-value below 0.05 is typically considered statistically significant, indicating that there is less than a 5% probability that the observed correlation occurred by chance. Thus, correlations with $p<0.05$ can be interpreted as significant at the 95% confidence level.
@@
CE is a statistical measure used to evaluate the predictive skill of a model by comparing its performance to that of a validation climatology [burger_verification_2007], thus indicating the shared variance between the observed and estimated data throughout the verification period. Given $n$ observations, where $y_{i}$ represents the actual values, $\hat{y}_{i}$ the predicted values, and $\bar{y}$ the mean of observed values, the coefficient of efficiency (CE) is computed for each grid point as,
${\mathrm{CE}}=1-\frac{\sum_{i=1}^{n}(y_{i}-\hat{y})^{2}}{\sum_{i=1}^{N}(y_{i}-\bar{y})^{2}}\penalty 10000\ .$ (5)
A value of $\rm{CE}=1$ indicates a perfect prediction, while $\rm{CE}=0$ suggests that the model performs no better than the validation climatology. Negative values imply that the model performs worse than the reference. In this study, we use the climatology of the 2011 – 2024 testing period as validation climatology. When the CE is evaluated during a specific season, only the related months are used to compute the validation climatology.
@@
To evaluate the aforementioned deterministic metrics on SEAS5, unless otherwise specified, we use the ensemble mean as the predicted value.
Continuous Ranked Probability Score (CRPS)
CRPS is a widely used proper scoring rule for ensemble predictions [Hersbach2020_crps]. The continuous ranked probability score (CRPS) measures the integrated squared difference between the cumulative distribution function (CDF) of the forecast ensemble, $F$, and the empirical CDF of the verifying observation, $H(\hat{y})=\mathbf{1}_{\{\hat{y}\geq y\}}$, where $y$ is the observed value. Formally, the CRPS is defined as
$\mathrm{CRPS}(F,y)=\int_{-\infty}^{\infty}\left[F(\hat{y})-H(\hat{y}-y)\right]^{2}\,d\hat{y}$, (6)
which reduces to the mean absolute error (MAE) in the case of deterministic forecasts. Unlike the MAE, the CRPS accounts for the full ensemble distribution. Lower CRPS values indicate higher forecast skill, with a score of zero corresponding to a perfect forecast.
In practice, when the forecast distribution is represented by a finite ensemble $\{f_{1},f_{2},\ldots,f_{m}\}$ of size $m$, the CRPS can be approximated by
$\mathrm{CRPS}(F,y)\approx\frac{1}{m}\sum_{i=1}^{m}\left|f_{i}-y\right|-\frac{1}{2m^{2}}\sum_{i=1}^{m}\sum_{j=1}^{m}\left|f_{i}-f_{j}\right|.$ (7)
The first term represents the mean absolute error of the ensemble members with respect to the observation, while the second term is a correction for the internal spread of the ensemble.
Spread–Skill Ratio (SSR)
The relationship between ensemble spread and forecast error is commonly assessed using the spread–skill ratio (SSR). For an ensemble of size $m$ with mean forecast $\bar{\hat{y}}_{t}$ at time $t$ and verifying observation $y_{t}$, the ensemble spread is defined as
$\sigma_{t}=\sqrt{\frac{1}{m-1}\sum_{i=1}^{m}\left(\hat{y}_{i,t}-\bar{\hat{y}}_{t}\right)^{2}},$ (8)
while the forecast error of the ensemble mean is measured by the root-mean-square error (RMSE),
$\mathrm{RMSE}=\sqrt{\frac{1}{T}\sum_{t=1}^{T}\left(\bar{\hat{y}}_{t}-y_{t}\right)^{2}}.$ (9)
The SSR is then given by the ratio between the temporal mean of the spread and the RMSE,
$\mathrm{SSR}=\frac{\overline{\sigma_{t}}}{\mathrm{RMSE}}.$ (10)
An ideal ensemble forecast is characterized by $\mathrm{SSR}\approx 1$, indicating that the predicted spread reliably reflects the actual forecast uncertainty. Values of $\mathrm{SSR}<1$ denote an under-dispersive and overconfident ensemble, in which the spread underestimates the true forecast error, whereas $\mathrm{SSR}>1$ indicates over-dispersion, with the ensemble spread being too large relative to the error.
@@ 5 Results
In the following sections, we evaluate the performance of the AI models in reconstructing monthly mean two-meter temperature and total precipitation anomalies in winter (DJF) and summer (JJA) from January 2011 to December 2024 . Sec. 5.1 presents the spatial reconstruction of these anomalies by the AI models trained with the seven year-round WR indices. We compare the performance of the AI model with the most commonly used reconstruction methods in Sec. 5.2. We further investigate the skill of the models with respect to: (a) the number of WR indices used for training and as input (Sec. 5.3); (b) the impact of errors on $I_{\rm wr}$ on the reconstruction (Sec. 5.4); and (c) the use of $I_{\rm wr}^{\rm SEAS5}$ as input (Sec. 5.5). The first two aspects address how the reconstruction quality degrades, identifying a threshold to achieve a better seasonal forecast than SEAS5. The third point examines the accuracy of the reconstructed anomalies when using WR indices predicted by SEAS5. It can be interpreted as a hybrid NWP-AI framework for forecasting European mean two-meter temperature and total precipitation in winter and summer.
Although the AI models reconstruct the anomalies over the entire European region as defined in Sec. 3.3, in the following analysis, we restrict the evaluation to land areas only.
@@ 5.1 Anomaly reconstruction with seven WR
Figure 3: Qualitative comparison of reconstructed monthly mean two-meter temperature (left column) and total precipitation (right column) against ERA5. The top row displays the cases with the lowest mean squared error, while the bottom row shows those with the highest. The best (worst) reconstruction for temperature corresponds to July 2016 (February 2011), while for precipitation it is August 2013 (November 2019).
A qualitative comparison between the reconstructed monthly mean two-meter temperature (left column) and total precipitation (right column) against ERA5 is shown in Fig. 3. The first row presents the best reconstruction, while the second row illustrates the worst reconstruction. Best and worst reconstructions are identified by the minimum and maximum MSE on the test set, respectively, evaluated after training. In the best-case scenario, the AI model is able to capture part of the spatial distribution of monthly two-meter temperature and total precipitation anomalies. However, it generally fails to reproduce extremes and more localized patterns, resulting in reduced spatial variability compared to the reanalysis. The worst-case scenarios usually occur when ERA5 anomalies are particularly strong. In such cases, the MSE increases if the model does not adequately reproduce either the magnitude or the spatial distribution of the anomalies. A representative example is the monthly total precipitation: while the model successfully mimics the spatial distribution in parts of Central Europe, as well as the Iberian and Scandinavian Peninsulas, it underestimates the magnitude of the extremes, which in this case reach up to 20 cm. The underestimation of extremes is a common issue for many NN architectures trained using the MSE.
Figure 4: AI models spatial MAE (top), ACC (middle) and CE (bottom) in reconstructing the winters (first and third columns) and summers (second and fourth columns) mean two-meter temperature and total precipitation anomalies using seven WR indices in the test years (from January 2011 to 2024 . Hatched areas mark regions with correlations that are not significant at $p>0.05$. Black contours indicate regions with ACC (CE) higher than 0.5 (0.3).
@@
Fig. 4 shows the spatial MAE, ACC and CE in reconstructing the European monthly mean two-meter temperature and total precipitation anomaly using the seven year-round WR indices, $I_{\rm wr}$, in the winters and summers during the test years.
@@
For the two-meter temperature anomaly, the MAE maps indicate low reconstruction errors across most of Europe, especially in western, central, and southern regions, with values frequently below 1.5K. Errors are slightly higher in northern and eastern areas, particularly during winter, which is related to the higher variability and amplitude of the two-meter temperature anomalies in those regions, as demonstrated by the higher standard deviation (not shown). The ACC maps show high correlations (generally above 0.5) between reconstructed and observed anomalies, particularly over western, central, and southern Europe in both winter and summer, as well as in some parts of Northern Europe. This indicates the model effectively captures temporal variability in these areas. The CE metric partially mirrors this pattern, with values approaching 0.3 in winter, signifying that the model performs appreciably better than a climatological reference. In summer, the CE remains low but generally above zero.
@@
In the case of precipitation, the MAE values are more spatially variable and generally higher than for temperature. Nonetheless, several regions, especially parts of Eastern and Northern Europe, exhibit relatively low MAE, particularly in winter. The ACC values for precipitation in winter are generally above 0.5 across most regions of Europe, except for some areas in central and eastern Europe. In the summer season, the ACC values remain positive but are generally low, with large regions falling below 0.5 and correlations that are not statistically significant ($p>0.05$). During this season, the AI model driven by the WR indices is largely unable to capture precipitation trends accurately, except in some regions of northern Europe. CE values for precipitation exhibit a similar spatial pattern of ACC, with positive skill in central Europe and part of Scandinavia during winter but a marked decline in the central and eastern regions during summer.
@@
Summer extratropical variability is less influenced by the large amplitude stationary waves and tropical forcing that characterize the boreal winter climate. As a result, the summer weather in the Northern Hemisphere is more influenced by regional-scale processes, such as mesoscale convective weather systems and land-atmosphere coupling [ambrizzi_rossby_1995, Randal_2000, Siegfried_2011]. This explains the lower skill reached by the AI models in summer compared to winter in all the skills tested. In particular, summer CE reveals that models using the Euro-Atlantic WR are unable to accurately reconstruct the mean two-meter temperature and total precipitation anomalies in the Mediterranean and Black Sea regions. For these areas, the reconstruction could be improved by incorporating the Mediterranean large-scale circulation [mastrantonas_extreme_2021] and including other predictors, such as soil moisture or rainfall deficit indices [vautard_summertime_2007]. In fact, the authors of the latter study demonstrate that “hot summers are preceded by winter rainfall deficits over Southern Europe”. Including indices that capture rainfall deficits (e.g., the Standardized Precipitation Index, SPI) and/or soil moisture (e.g., the Soil Moisture Deficit Index, SMDI) from the preceding season could enhance the reconstruction of the Southern European climate. We hypothesize that our model would be capable of capturing the relationship between these indices and the monthly summer anomalies of two-meter temperature and total precipitation in the region.
5.2 Comparison with regressions on individual WR
We compare the performance of the AI models against a benchmark based on monthly composites of temperature and precipitation for each of the seven WR [see, e.g., grams_balancing_2017, madonna_reconstructing_2021, gerighausen_quantifying_2024]. We first identify the dominant WR for each month in the winters and summers from 1981 to 2010. A WR is considered dominant if its monthly index exceeds all other WR indices and is larger than its standard deviation over the climatology. We then compute the typical monthly two-meter temperature and total precipitation associated with each WR by averaging the respective values over all months in which that WR was dominant. Months without a clearly dominant WR are assigned to a “no-regime” composite. Since each WR can have a different impact on the ground variables depending on the season, the average is computed separately for the months in the winter (DJF) and summer (JJA) season. For the reconstruction period (2011 onward), we assign to each month the composite corresponding to its active dominant WR, or the no-regime composite if no dominant WR could be identified, applying the same classification used in the climatology. As for the AI model, this approach relies on monthly average indices as predictors, enabling a direct comparison.
The composite reconstructions are evaluated using the deterministic metrics introduced in 4.2 against ERA5 monthly data. Relative scores are computed with respect to the AI-model as:
  • relative improvement of the MAE: (Composite MAE - AI-Model MAE) / Composite MAE,
  • relative improvement of the ACC: (AI-Model ACC - Composite ACC) / (1 - Composite ACC),
  • relative improvement of the CE: (AI-Model CE - Composite CE) / (1 - Composite CE).
These scores highlight the relative improvement of the AI model with respect to the regressions on individual WR as a benchmark.
Figure 5: Relative improvement of the AI models with respect to the regressions on individual WR as benchmark. First (second) [third] row shows the relative improvement in MAE (ACC) [CE] for both two-meter temperature and total precipitation separately for the winter (December, January, February) and summer (June, July, August) from 2011 to 2024 . Solid (dashed) contours indicate regions with relative improvement higher (lower) than 30%.
Fig. 5 shows the relative improvement in percentage for both the two-meter temperature and total precipitation reconstruction in winter and summer from 2011 to 2024 . The AI models outperform the simple composite benchmark across all seasons, variables, and metrics, with only a few regional exceptions. The improvement is particularly relevant for ACC and CE of the two-meter temperature reconstruction in central and Mediterranean Europe for both seasons, and particularly during summer, with the exception of some areas in Norway, Sweden, Greenland, and the British Isles. Also, the ACC and CE of the total precipitation show substantial improvements in many areas, but with less spatial consistency and with few regions characterized by negative scores in the Iberian peninsula.
Overall, the comparison demonstrates that while the composite method provides a simple and interpretable benchmark, the AI models are capable of capturing subtle combinations of WR patterns and reproducing regional variability more accurately. These results highlight the added value of data-driven reconstructions over traditional composite approaches, particularly when the underlying regime dynamics are complex.